OpenStudy (anonymous):

Perform the indicated operation and, if possible simplify. 10/33 * (-5) * (-11/24).... thanks for your help in advance.

OpenStudy (anonymous):

=-55/33 *(-11/24)=55/3 *1/24=55/72

jhonyy9 (jhonyy9):

why -55/33 --- this not is right because 10 multiplied with -5 is -50 and not is -55 will be : (-50/33 )(-11/24) = 550/792 =275/396

OpenStudy (anonymous):

\[(10/33)*(-5)*(-11/24)=\frac{25}{36} \]

OpenStudy (anonymous):

\[275/396=\frac{25}{36} \]